# Generalized invariants meet constitutive neural networks: A novel framework for hyperelastic materials

### v2

@@ 1 Introduction
The mechanical behavior of hyperelastic solids is typically encoded in a strain energy function $\psi$ whose derivatives can be used to map the deformation gradient onto the local stress state. For isotropic incompressible hyperelastic materials, classical models commonly express the strain energy function $\psi$ in terms of the first and second invariants $I_{1}$ and $I_{2}$ of the right Cauchy–Green deformation tensor. Several commonly used isotropic hyperelastic models rely on different combinations of these invariants. The Neo-Hooke model depends solely on $I_{1}$ [27], the Blatz-Ko model on $I_{2}$ [4], the Mooney–Rivlin model combines $I_{1}$ and $I_{2}$ [24], and the Yeoh model incorporates only $I_{1}$, but uses it in a higher-order polynomial expansion [34]. While these models offer analytical convenience and they are useful for moderate deformations, they often generalize poorly to multiaxial or shear-dominated loading regimes, particularly at large strains [8, 20].
To improve the expressivity and predictive power of constitutive models, recent advances have focused on data-driven approaches, such as neural networks trained to approximate the strain energy function directly from experimental observations. These strategies include constitutive artificial neural networks (CANNs) [14, 15, 16], physics-informed neural networks (PINNs) [26, 9], or physics-augmented neural networks (PANNs) [13] and encode physical constraints such as material frame indifference throughout the careful selection of invariant-based inputs and network architecture. In principal-stretch-based models, neural networks use principal stretches as input [29, 30]. This formulation allows for a high degree of flexibility and has proven effective in modeling soft tissues, where mechanical responses are strongly non-linear. Very recently, material fingerprinting, a novel technology that abandons the use of neural networks and avoids solving non-linear optimization problems, has been proposed as an effective alternative model discovery [7].
@@ 2.1.1 Two-step approach
$\frac{\sigma_{11}}{\sigma_{22}}=\frac{\alpha\lambda_{1}^{\alpha}-\alpha(\lambda_{1}\lambda_{2})^{-\alpha}}{\alpha\lambda_{2}^{\alpha}-\alpha(\lambda_{1}\lambda_{2})^{-\alpha}}.$ (8)
This expression is entirely independent of the form of the strain energy function, making it suitable to directly estimate the power $\alpha$. For a dataset of measured stress components $\hat{\sigma}_{11},\hat{\sigma}_{22}\,$} and corresponding stretches $\lambda_{1},\lambda_{2}\,$}, the loss function is defined as
@@
where $\mu>0$ is a stress-like material constant, and the one-term Anssari‑Benam Bucchi strain energy [1],
@@
where $n>0$ and $N>0$ are additional dimensionless model parameters. Notably, when $N\to\infty$, we recover the one-term Ogden model in (10). Given a fixed $\alpha^{*}$ from the first step, we identify the model parameters $\mathbf{w}=\{\,\mu\,\}$ or $\mathbf{w}=\{\,n,N\,\}$, by minimizing the loss function that compares the predicted stress to the measured data,
$L_{2}(\mathbf{w};\mathcal{J}_{\alpha^{*}}(\mathbf{F}))=\frac{1}{n_{\text{data}}}\sum_{i=1}^{n_{\text{data}}}\left\|\,\bm{\sigma}^{\alpha^{*}}(\mathcal{J}_{\alpha^{*}}(\mathbf{F}_{i}),\mathbf{w})-\hat{\bm{\sigma}}_{i}\,\right\|_{2}^{2}\;\rightarrow\;\min.$ (12)
where $\bm{\sigma}^{\alpha^{*}}$ are the modeled Cauchy stresses according to (5) for a specific strain energy function $\psi^{\mathrm{OG}}$ in (10) or $\psi^{\mathrm{AB}}$ in (11).
@@ 2.2 Constitutive neural networks for model discovery
$L(\alpha,\mathbf{w};\mathbf{F})=\frac{1}{n_{\text{data}}}\sum_{i=1}^{n_{\text{data}}}\left\|\,\bm{\sigma}^{\alpha}(\mathcal{J}_{\alpha}(\mathbf{F}_{i}),\mathbf{w})-\hat{\bm{\sigma}}_{i}^{\alpha}\,\right\|_{2}^{2}+r\,\sum_{i=1}^{n_{\rm{weights}}}|\,w_{i}\,|\;\rightarrow\;\min.$ (14)
@@
PI-CANN [29] a priori given
@@
Neo-Hooke $\mathrm{w}_{1}^{\mathrm{id}}[\mathcal{J}_{2}-3]$ $\{2\}$ $\{1\}$ id
@@ 2.3 Benchmarking
To evaluate model performance, we compare the computationally predicted and experimentally measured responses using the coefficient of determination $R^{2}$, which quantifies the goodness of fit. We assess and compare the goodness of fit across various approaches, including SI-CANN, PI-CANN and our proposed GI-CANN as well as results from [6] comparing unconstrained feedforward neural networks (FNN) and fully input convex neural networks (FICNN) using the same training and testing data. We minimize the loss function (14) using the Adam algorithm [10].
@@ 3 Results and Discussion
combines an identity function and an exponential term, both formulated on the same generalized invariant $\mathcal{J}_{\alpha_{+}^{*}}$ with $\alpha_{+}^{*}=1.675$. This formulation results in an excellent fit across all deformation modes, achieving an average coefficient of determination $R^{2}=0.998$, as detailed in Table 3. The visual evidence of this accuracy is confirmed by the first row of Figure 2, particularly in columns three through six, where the modeled responses closely align with the experimental data shown as white circles. Red-shaded regions dominate the visualizations, indicating a strong influence of positive $\alpha_{+}$ terms on the overall Cauchy stress components. These regions highlight the model’s reliance on $I_{1}$-like effects. We can further improve the model predictions by including both positive and negative exponents $\alpha$. In these cases, the regularization embedded in the GI-CANN architecture selectively activates only those components that contribute meaningfully to the stress response. This selective activation is particularly evident in models that use the combined invariant set $\{\mathcal{J}_{\alpha_{+}},\mathcal{J}_{\alpha_{-}}\}$. The two-invariant model achieves a near-perfect fit $R^{2}=0.999$, as shown in the bottom section of Table 3, demonstrating the advantage of incorporating flexible representations of the strain energy function. In contrast, models that use only negative exponents $\alpha_{-}$, intended to reflect $I_{2}$-like behavior, perform poorly for rubber. These models fail to capture the dominant stretch-induced behavior and yield low accuracy, with average $R^{2}$ values as low as 0.195 for the FICNN and 0.264 for the GI-CANN. The corresponding visualizations show faint or absent blue regions, confirming the minimal contribution of negative $\alpha$ terms to the stress response.
@@
For the brain data, a different mechanical behavior emerges. Human brain cortex, which deforms under much smaller strains and exhibits shear-dominant and pressure-sensitive characteristics, responds best to generalized invariants with large negative exponent values. These values correspond to an $I_{2}$-like behavior. Remarkably, even a single invariant $\mathcal{J}_{\alpha_{-}^{*}}$ with an optimally chosen negative exponent $\alpha_{-}^{*}=-18.572$ proves sufficient to accurately describe the tissue’s response under all loading conditions, uniaxial tension, uniaxial compression, and previously unseen simple shear. The resulting one-term strain energy function,
$\psi=\mathrm{f_{id}}(\mathcal{J}_{-18.572})=0.009\,\mathrm{kPa}\cdot(\mathcal{J}_{-18.572}-3),$ (20)
@@
produces an average $R^{2}=0.967$, significantly surpassing traditional models built around $I_{1}$ or $I_{2}$. As shown in Figure 3, the third row visualizes the dominance of negative $\alpha_{-}$ contributions through strong blue regions in the stress maps. These patterns emphasize the model’s reliance on $I_{2}$-like deformation measures and the negligible role of any positive $\alpha_{+}$ values, which are effectively pruned by the model’s sparsity-promoting regularization. Indeed, in the columns corresponding to $\alpha_{+}>0$, the red-shaded regions vanish, confirming the absence of positive $\alpha_{+}$ contributions. Among $2^{3}-1$ possible isotropic models based on a single generalized invariant, our automatically discovered model in (20) exactly aligns with findings in previous studies [5, 22, 23, 21] where the models have been manually selected. For example, among five traditional models, Neo-Hooke, Mooney-Rivlin, Demiray, Gent, and Ogden, the one-term Ogden model with a high negative exponent was identified as the only model that best represents the response of human brain tissue under different loading conditions [5]. Notably, the SI-CANN does not have the flexibility to discover the Ogden model for arbitrary exponent values, whereas the PI-CANN cannot identify models involving exponential or logarithmic functions. Our new GI-CANN inherently eliminates both these limitations. Due to the relatively small deformation during the brain experiments, the classical invariants $I_{1}$ and $I_{2}$ have nearly identical values. As a result, standard models based on either of these invariants produce similar fitting results. This behavior is reflected in the first and second rows of Figure 3, where the visualizations across the $I_{1}$- and $I_{2}$-like inputs appear nearly indistinguishable. Table 4 corroborates this finding, with models using $\alpha=2$ or $\alpha=-2$ both yielding moderate $R^{2}$ values, but falling short of the high fidelity achieved with the optimized negative-exponent model. This observation aligns well with previous findings [6] and underscores the power of generalized invariants, significantly improving the goodness of fit compared to standard-invariant-based models. Although the FNN model achieves the highest average $R^{2}=0.982$ when combining both positive and negative invariants with fixed exponents, this result does not translate into a structurally interpretable formulation. Instead, the GI-CANN model demonstrates nearly the same accuracy with $R^{2}=0.966$, while maintaining interpretability and sparsity. The network naturally suppresses less relevant inputs, a fact confirmed by the absence of the red-shaded stress contributions in the third row of Figure 3. Moreover, the PS-CANN model with a dense range of discrete exponents from 30-30 to $30$ also performs competitively, but lacks the efficiency and interpretability of GI-CANN’s learned invariant representations.
@@
Taken together, our findings corroborate prior insights reported in the literature. For example, previous suggested that optimal exponents in generalized-invariant-based models differ significantly between elastomers and biological tissues [3], but previous studies required a two-step fitting process and full stress–stretch datasets. Our GI-CANN achieves comparable and often superior results with single-step training and reduced data requirements. The model’s predictive power, evaluated by the coefficient of determination $R^{2}$, consistently outperforms standard invariant neural models, constitutive neural networks [15], unconstrained feedforward neural networks, and fully input convex neural network [6] across all test configurations.
@@
FICNN [6] $\mathcal{J}_{\alpha_{+}}$ $\alpha_{+}=2$ 0.987 0.966 0.968 0.973
FNN [6] $\mathcal{J}_{\alpha_{+}}$ $\alpha_{+}=2$ 0.987 0.966 0.968 0.978
@@
FNN [6] $\mathcal{J}_{\alpha_{-}}$ $\alpha_{-}=-2$ 0 0.790 0 0.459
@@
GI-CANN $\mathcal{J}_{\alpha_{-}}$ $\alpha^{*}_{-}=-18.572$ 0.930 0.997 0.974 0.967
@@ 4 Conclusion and Outlook
For rubber, our findings confirm that the optimal invariant structure is closely aligned with the classical first invariant $I_{1}$, consistent with well-established models including the Neo-Hooke or Mooney–Rivlin formulations, that reflect the general notion that rubber exhibits an entropic elasticity, meaning that the majority of its elastic response arises from changes in chain conformational entropy, rather than from changes in internal energy [32]. At the macrosocopic level, the first invariant captures the average chain extension [11] and the elastic energy in rubber is mostly stored through entropic effects related to molecular chain extension [32]. The best-performing GI-CANN models feature positive exponents $\alpha_{+}$, suggesting that stretch-dominant behaviors are well captured by $\mathcal{J}_{\alpha_{+}}$-type invariants. For brain tissue, in contrast, the data are better represented using $\mathcal{J}_{\alpha_{-}}$ with large negative exponents. This suggests that the molecular origin of elasticity in brain tissue is fundamentally different than in rubber.
@@
Despite these promising results, several limitations remain, opening the door for further investigations. First, our current implementation is restricted to isotropic materials and assumes incompressibility. Extending the GI-CANN to anisotropic and compressible formulations would require careful adaptation of the invariant structures and associated input features. Second, although the GI-CANN discovers optimal exponents $\alpha^{*}$ as part of its training, the physical interpretability of highly negative or fractional values remains an open question [2]. Third, the training process, while automated, still involves hyperparameter tuning that involves the neural network architecture and the regularization process, which may both benefit from further optimization. Moreover, while the GI-CANN performs well, even with limited experimental data, the accuracy of the discovered models is inherently tied to the quality and diversity of the training data. In particular, if the loading conditions fail to sufficiently excite the relevant deformation modes, the learned invariants may not generalize. This issue is especially pertinent for biological tissues, where experimental datasets are often sparse or noisy.
@@
Overall, generalized-invariant-based constitutive artificial neural networks provide a compelling tool to discover interpretable models for soft matter systems by combining physical insight with data-driven flexibility. Future work should explore extensions to time-dependent behavior, anisotropy, and uncertainty quantification to broaden the applicability, robustness, and adaptation of the approach.

### v3

@@ 1 Introduction
The mechanical behavior of hyperelastic solids is typically encoded in a strain energy function $\psi$ whose derivatives can be used to map the deformation gradient onto the local stress state. For isotropic incompressible hyperelastic materials, classical models commonly express the strain energy function $\psi$ in terms of the first and second invariants $I_{1}$ and $I_{2}$ of the right Cauchy–Green deformation tensor. Several commonly used isotropic hyperelastic models rely on different combinations of these invariants. The neo-Hookean model depends solely on $I_{1}$ [30], the Blatz-Ko model on $I_{2}$ [4], the Mooney–Rivlin model combines $I_{1}$ and $I_{2}$ [27], and the Yeoh model incorporates only $I_{1}$, but uses it in a higher-order polynomial expansion [37]. While these models offer analytical convenience and they are useful for moderate deformations, they often generalize poorly to multiaxial or shear-dominated loading regimes, particularly at large strains [10, 23].
To improve the expressivity and predictive power of constitutive models, recent advances have focused on data-driven approaches, such as neural networks trained to approximate the strain energy function directly from experimental observations. These strategies include constitutive artificial neural networks (CANNs) [16, 18, 19], physics-informed neural networks (PINNs) [29, 11], or physics-augmented neural networks (PANNs) [15] and encode physical constraints such as material frame indifference throughout the careful selection of invariant-based inputs and network architecture. In principal-stretch-based models, neural networks use principal stretches as input [32, 33]. This formulation allows for a high degree of flexibility and has proven effective in modeling soft tissues, where mechanical responses are strongly nonlinear. Very recently, material fingerprinting, a novel technology that abandons the use of neural networks and avoids solving nonlinear optimization problems, has been proposed as an effective alternative model discovery [8].
@@ 2.1.1 Two-step approach
$\frac{\sigma_{11}}{\sigma_{22}}=\frac{\lambda_{1}^{\alpha}-(\lambda_{1}\lambda_{2})^{-\alpha}}{\lambda_{2}^{\alpha}-(\lambda_{1}\lambda_{2})^{-\alpha}}.$ (8)
Alternatively, we could use the chain rule by first computing the derivatives of the generalized invariants with respect to the classical first and second invariants [3]. The expression in (8) is entirely independent of the form of the strain energy function, making it suitable to directly estimate the power $\alpha$. For a dataset of measured stress components $\hat{\sigma}_{11},\hat{\sigma}_{22}\,$} and corresponding stretches $\lambda_{1},\lambda_{2}\,$}, the loss function is defined as
@@
where $\mu>0$ is a stress-like material constant, and the one-term Anssari‑Benam strain energy [1],
@@
where $n>0$ and $N>0$ are additional dimensionless model parameters. Notably, when $N\to\infty$, we recover the one-term Ogden model in (10). Given a fixed $\alpha^{*}$ from the first step, we identify the model parameters $\mathbf{w}=\{\,\mu\,\}$ or $\mathbf{w}=\{\,n,N\,\}$, by minimizing the loss function,
$L_{2}(\mathbf{w};\mathcal{J}_{\alpha^{*}}(\mathbf{F}))=\frac{1}{n_{\text{data}}}\sum_{i=1}^{n_{\text{data}}}\left\|\,\bm{\sigma}^{\alpha}(\mathcal{J}_{\alpha^{*}}(\mathbf{F}_{i}),\mathbf{w})-\hat{\bm{\sigma}}_{i}\,\right\|_{2}^{2}\;\rightarrow\;\min.$ (12)
where $\bm{\sigma}^{\alpha}$ are the modeled Cauchy stresses according to (5) for a specific strain energy function $\psi^{\mathrm{OG}}$ in (10) or $\psi^{\mathrm{AB}}$ in (11) and $\hat{\bm{\sigma}}_{i}$ denote the measured data.
@@ 2.2 Constitutive neural networks for model discovery
$L(\alpha,\mathbf{w};\mathbf{F})=\frac{1}{n_{\text{data}}}\sum_{i=1}^{n_{\text{data}}}\left\|\,\bm{\sigma}^{\alpha}(\mathcal{J}_{\alpha}(\mathbf{F}_{i}),\mathbf{w})-\hat{\bm{\sigma}}_{i}\,\right\|_{2}^{2}+r\,\sum_{i=1}^{n_{\rm{weights}}}|\,w_{i}\,|\;\rightarrow\;\min,$ (14)
@@
PS-CANN [32] a priori given
@@
neo-Hookean $\mathrm{w}_{1}^{\mathrm{id}}[\mathcal{J}_{2}-3]$ $\{2\}$ $\{1\}$ id
@@ 2.3 Benchmarking
To evaluate model performance, we compare the computationally predicted and experimentally measured responses using the coefficient of determination $R^{2}$, which quantifies the goodness of fit. We assess and compare the goodness of fit across various approaches, including SI-CANN, PS-CANN and our proposed GI-CANN as well as results from [6] comparing unconstrained feedforward neural networks (FNN) and fully input convex neural networks (FICNN) using the same training and testing data. We minimize the loss function (14) using the Adam algorithm [12].
@@ 3 Results and Discussion
combines an identity function and an exponential term, both formulated on the same generalized invariant $\mathcal{J}_{\alpha_{+}^{*}}$ with $\alpha_{+}^{*}=1.675$. This formulation results in an excellent fit across all deformation modes, achieving an average coefficient of determination $R^{2}=0.998$, as detailed in Table 3. The visual evidence of this accuracy is confirmed by the first row of Figure 2, particularly in columns three through six, where the modeled responses closely align with the experimental data shown as white circles. Red-shaded regions dominate the visualizations, indicating a strong influence of positive $\alpha_{+}$ terms on the overall Cauchy stress components. These regions highlight the model’s reliance on $I_{1}$-like effects. The relative error plots in Figure 4 further confirm the accuracy of the discovered model in (19), which outperforms the discovered standard-invariant-based model, $\psi(I_{1})=\rm{f_{id}}(I_{1})+\rm{f_{exp}}(I_{1})$, with the best performance achieved under uniaxial tension. Compared to the error plots of the Anssari-Benam model (11), our discovered model produces smaller errors at low stretches but larger errors at high stretches. Nevertheless, the strength of the proposed approach lies in its ability to automatically identify a suitable model rather than relying on a priori assumptions about the strain energy function. We can further improve the model predictions by including both positive and negative exponents $\alpha$. In these cases, the regularization embedded in the GI-CANN architecture selectively activates only those components that contribute meaningfully to the stress response. This selective activation is particularly evident in models that use the combined invariant set $\{\mathcal{J}_{\alpha_{+}},\mathcal{J}_{\alpha_{-}}\}$. The two-invariant model achieves a near-perfect fit $R^{2}=0.999$, as shown in the bottom section of Table 3, demonstrating the advantage of incorporating flexible representations of the strain energy function. In contrast, models that use only negative exponents $\alpha_{-}$, intended to reflect $I_{2}$-like behavior, perform poorly for rubber. These models fail to capture the dominant stretch-induced behavior and yield low accuracy, with average $R^{2}$ values as low as 0.195 for the FICNN and 0.264 for the GI-CANN. The corresponding visualizations show faint or absent blue regions, confirming the minimal contribution of negative $\alpha$ terms to the stress response.
@@
For the brain data, a different mechanical behavior emerges. Human brain cortex, which deforms under much smaller strains and exhibits shear-dominant and pressure-sensitive characteristics, responds best to generalized invariants with large negative exponent values. These values correspond to an $I_{2}$-like behavior. Remarkably, even a single invariant $\mathcal{J}_{\alpha_{-}^{*}}$ with an optimally chosen negative exponent $\alpha_{-}^{*}=-18.016$ proves sufficient to accurately describe the tissue’s response under all loading conditions, uniaxial tension, uniaxial compression, and previously unseen simple shear. The resulting one-term strain energy function,
$\psi=\mathrm{f_{id}}(\mathcal{J}_{-18.016})=0.009\,\mathrm{kPa}\cdot(\mathcal{J}_{-18.016}-3),$ (20)
@@
produces an average $R^{2}=0.966$, significantly surpassing traditional models built around $I_{1}$ or $I_{2}$. As shown in Figure 3, the third row visualizes the dominance of negative $\alpha_{-}$ contributions through strong blue regions in the stress maps. These patterns emphasize the model’s reliance on $I_{2}$-like deformation measures and the negligible role of any positiv $\alpha_{+}$ values, which are effectively pruned by the model’s sparsity-promoting regularization. Indeed, in the columns corresponding to $\alpha_{+}>0$, the red-shaded regions vanish, confirming the absence of positive $\alpha_{+}$ contributions. Among $2^{3}-1$ possible isotropic models based on a single generalized invariant, our automatically discovered model in (20) exactly aligns with findings in previous studies [5, 25, 26, 24] where the models have been manually selected. For example, among five traditional models, neo-Hookean, Mooney-Rivlin, Demiray, Gent, and Ogden, the one-term Ogden model with a high negative exponent was identified as the only model that best represents the response of human brain tissue under different loading conditions [5]. Moreover, the identified model parameters are close to the parameters recently identified via the efficient unsupervised constitutive law identification EUCLID [7, 9]. Notably, the SI-CANN does not have the flexibility to discover the Ogden model for arbitrary exponent values, whereas the PS-CANN cannot identify models involving exponential or logarithmic functions. Our new GI-CANN inherently eliminates both these limitations. The relative error plots in Figure 5 show that the discovered model in (20) again produces significantly smaller errors than the discovered standard-invariant-based model $\psi(I_{2})=\rm{f_{ln}}(I_{2})$, which corresponds to the Gent model. Despite the small-strain regime and the complexity of soft tissue mechanics, these results slightly outperform, or at least match, those of the Anssari-Benam model [1]. The strength of the proposed method is again evident: It automatically discovers a well-suited model without a priori assumptions about the strain energy function and naturally adapts to the tissue’s shear-dominant and pressure-sensitive response.
Due to the relatively small deformation during the brain experiments, the classical invariants $I_{1}$ and $I_{2}$ have nearly identical values. As a result, standard models based on either of these invariants produce similar fitting results. This behavior is reflected in the first and second rows of Figure 3, where the visualizations across the $I_{1}$- and $I_{2}$-like inputs appear nearly indistinguishable. Table 4 corroborates this finding, with models using $\alpha=2$ or $\alpha=-2$ both yielding moderate $R^{2}$ values, but falling short of the high fidelity achieved with the optimized negative-exponent model. This observation aligns well with previous findings [6] and underscores the power of generalized invariants, significantly improving the goodness of fit compared to standard-invariant-based models. Although the FNN model achieves the highest average $R^{2}=0.982$ when combining both positive and negative invariants with fixed exponents, this result does not translate into a structurally interpretable formulation. Instead, the GI-CANN model demonstrates nearly the same accuracy with $R^{2}=0.966$, while maintaining interpretability and sparsity. The network naturally suppresses less relevant inputs, a fact confirmed by the absence of the red-shaded stress contributions in the third row of Figure 3. Moreover, the PS-CANN model with a dense range of discrete exponents from 30-30 to $30$ also performs competitively, but lacks the efficiency and interpretability of GI-CANN’s learned invariant representations.
@@
Taken together, our findings corroborate prior insights reported in the literature. For example, previous work suggested that the optimal exponents in generalized-invariant-based models differ noticeably across different materials [3], but those studies required a two-step fitting process and full stress–stretch datasets. Our GI-CANN achieves comparable and often superior results with single-step training and reduced data requirements. The model’s predictive power, evaluated by the coefficient of determination $R^{2}$, consistently outperforms standard invariant neural models, constitutive neural networks [18], unconstrained feedforward neural networks, and fully input convex neural network [6] across all test configurations.
@@
FICNN [6] $\mathcal{J}_{\alpha_{+}}$ $\alpha_{+}=2$ 0.987 0.966 0.954 0.969
FNN [6] $\mathcal{J}_{\alpha_{+}}$ $\alpha_{+}=2$ 0.983 0.984 0.968 0.978
@@
FNN [6] $\mathcal{J}_{\alpha_{-}}$ $\alpha_{-}=-2$ 0 0.790 0.589 0.460
@@
Figure 4: Relative error plots for rubber models. Comparison between the generalized-invariant-based model $\psi(\mathcal{J}_{\alpha})$ in (19) and a model $\psi(I_{1})$ based on the standard invariant $I_{1}$. The relative error is computed as $\left|\frac{\sigma_{ij}^{\alpha}-\hat{\sigma}_{ij}}{\hat{\sigma}_{ij}}\right|\times 100$, where $\sigma_{ij}^{\alpha}$ and $\hat{\sigma}_{ij}$ denote the modeled and measured Cauchy stress components, respectively. From left to right: uniaxial tension (ut), equibiaxial tension (bt), and pure shear (ps).
@@
GI-CANN $\mathcal{J}_{\alpha_{-}}$ $\alpha^{*}_{-}=-18.016$ 0.930 0.998 0.970 0.966
@@
Figure 5: Relative error plots for brain cortex models. Comparison between the generalized-invariant-based model $\psi(\mathcal{J}_{\alpha})$ in (20) and a model $\psi(I_{2})$ based on the standard invariant $I_{2}$. The relative error is computed as $\left|\frac{\sigma_{ij}^{\alpha}-\hat{\sigma}_{ij}}{\hat{\sigma}_{ij}}\right|\times 100$, where $\sigma_{ij}^{\alpha}$ and $\hat{\sigma}_{ij}$ denote the modeled and measured Cauchy stress components, respectively. From left to right: uniaxial tension (ut), uniaxial compression (uc), and simple shear ($s$).
@@ 4 Conclusion and Outlook
For rubber, our findings confirm that the optimal invariant structure is closely aligned with the classical first invariant $I_{1}$, consistent with well-established models including the neo-Hookean or Mooney–Rivlin formulations, that reflect the general notion that rubber exhibits an entropic elasticity, meaning that the majority of its elastic response arises from changes in chain conformational entropy, rather than from changes in internal energy [35]. At the macrosocopic level, the first invariant captures the average chain extension [13] and the elastic energy in rubber is mostly stored through entropic effects related to molecular chain extension [35]. The best-performing GI-CANN models feature positive exponents $\alpha_{+}$, suggesting that stretch-dominant behaviors are well captured by $\mathcal{J}_{\alpha_{+}}$-type invariants. For brain tissue, in contrast, the data are better represented using $\mathcal{J}_{\alpha_{-}}$ with large negative exponents. This suggests that the molecular origin of elasticity in brain tissue is fundamentally different than in rubber.
@@
Despite these promising results, several limitations remain, opening the door for further investigations. First, our current implementation is restricted to isotropic materials and assumes incompressibility. Extending the GI-CANN to anisotropic or compressible formulations would require careful adaptation of the invariant structures and associated input features. Second, although the GI-CANN discovers optimal exponents $\alpha^{*}$ during training, the physical interpretation of highly negative values can be challenging [2]. However, similar exponents have been reported in previous studies [5, 25, 26, 24, 9], indicating that such values are commonly used to capture nonlinear material behavior. While these exponents improve predictive accuracy, their mechanistic meaning is still partially open, and future work could incorporate physics-informed constraints [29] or Bayesian priors [17] to guide $\alpha$ toward more interpretable ranges. Third, the training process, while automated, still requires hyperparameter tuning of both the network architecture and the regularization process. The current choices were selected empirically based on validation performance and motivated by previous works to ensure comparability. Finally, while GI-CANN performs well even with limited experimental data, the accuracy of discovered models depends on the quality and diversity of the training data. In particular, if the loading conditions fail to sufficiently excite the relevant deformation modes, the learned invariants may not generalize. This issue is especially pertinent for biological tissues, where experimental datasets are often sparse or noisy. However, at present, no universal protocol exists, but we chose the deformation modes reported here to ensure comparability with prior studies [18, 19, 6].
@@
Overall, generalized-invariant-based constitutive artificial neural networks provide a compelling tool to discover interpretable models for soft matter systems by combining physical insight with data-driven flexibility. Future work will explore extensions to time-dependent behavior, anisotropy, and uncertainty quantification to broaden the applicability, robustness, and adaptation of the approach.